# Decomposing Multivariate Information Rates in Networks of Random Processes

### v2

@@ Abstract
The Partial Information Decomposition (PID) framework has emerged as a powerful tool for analyzing high-order interdependencies in complex network systems. However, its application to dynamic processes remains challenging due to the implicit assumption of memorylessness, which often falls in real-world scenarios. In this work, we introduce the framework of Partial Information Rate Decomposition (PIRD) that extends PID to random processes with temporal correlations. By leveraging mutual information rate (MIR) instead of mutual information (MI), our approach decomposes the dynamic information shared by multivariate random processes into unique, redundant, and synergistic contributions obtained aggregating information rate atoms in a principled manner. To solve PIRD, we define a pointwise redundancy rate function based on the minimum MI principle applied locally in the frequency-domain representation of the processes. The framework is validated in benchmark simulations of Gaussian systems, demonstrating its advantages over traditional PID in capturing temporal correlations and showing how the spectral representation may reveal scale-specific higher-order interactions that are obscured in the time domain. Furthermore, we apply PIRD to a physiological network comprising cerebrovascular and cardiovascular variables, revealing frequency-dependent redundant information exchange during a protocol of postural stress. Our results highlight the necessity of accounting for the full temporal statistical structure and spectral content of vector random processes to meaningfully perform information decomposition in network systems with dynamic behavior such as those typically encountered in neuroscience and physiology.
@@ I Introduction
Thanks to the peculiarities described above, the PID framework has become widely adopted as a main tool to assess high-order interdependencies among the units of network datasets collected in several applicative fields of physics, engineering and life sciences where understanding the specific contributions of individual variables to the collective information is of fundamental importance Wibral et al. (2017); Cang and Nie (2020); Rosas et al. (2020); Luppi et al. (2022); Wollstadt et al. (2023); Dissanayake et al. (2024). For instance, it has been used to analyze neural information processing dynamics, revealing how different brain regions encode stimuli both redundantly and synergistically Kay et al. (2022). In machine learning, PID has been exploited in feature selection to distinguish between information that is redundant across features and information that is uniquely informative Barnett et al. (2009a); Taylor (2018); Wollstadt et al. (2023). Recent applications have highlighted its value for network analysis also in computational neuroscience, biology and physiology Luppi et al. (2024a); Sherrill et al. (2021); Faes et al. (2021).
@@
Another crucial issue, which is the main question addressed in this paper, is how to apply PID to random processes with temporal statistical structure. In fact, although specifically defined for random variables, PID is often needed in practice to analyze multivariate time series whose most proper statistical representation is the vector random process. Then, since both the target random process and the set of source processes constitute collections of random variables, utilization of the PID in such dynamic case is not straightforward as it implies an arbitrary selection of the variables to be extracted from each process. In the literature, PID has been applied to the variables sampling the processes at the same time, performing a so-called "static PID", under the implicit assumption that the processes are stationary and memoryless (i.e., composed by independent and identically distributed - i.i.d. - variables). However, the i.i.d. assumption is typically not tested in practice and is often violated in applications of information decomposition where the analyzed data exhibit nontrivial temporal correlations Kay et al. (2022); Varley et al. (2023a, b). Alternatively, the PID has been applied to random processes by selecting the variables to use in order to decompose the joint information transferred from all sources to the target Krohova et al. (2019); Luppi et al. (2022); Koçillari et al. (2023); Luppi et al. (2024b). Specifically, taking the present of one predefined process as the target variable and the past histories of all other processes as (vector) source variables, and conditioning the MI between target and sources on the past of the predefined process, yields the PID of the popular transfer entropy (TE) measure of information transfer Schreiber (2000). Nevertheless, although considering the temporal statistical structure of the multivariate process, the PID of the TE cannot account for instantaneous interactions among the processes, nor for causal interactions occurring in the causal direction from the target to the sources. Hence, the current applications of PID to random processes provide only a partial, often misleading view of the interactions among the units of dynamic network systems.
@@
To deal with these significant limitations, the present work introduces a framework for the decomposition of the information shared dynamically between the target and the source units composing the analyzed network of random processes. Our idea is to shift the paradigm of PID from the use of random variables to the use of random processes as building blocks of information decomposition: leveraging the use of information rate quantities in place of standard information quantities, we replace the MI between random variables with the MI (MIR) between random processes, and use the same lattice backbone of PID to implement the so-called partial information rate decomposition (PIRD). PIRD dissects the information shared per unit of time between the designed target random process and the set of source processes, quantified by the MIR, into PI rate atoms adopting a full PID perspective, or into unique, redundant and synergistic information rates adopting a coarse-graining perspective. The PIRD is solved introducing the new information-theoretic measure of redundancy rate which generalizes the MIR over the lattice. Among several possible definitions, the redundancy rate is here formulated following a pointwise approach implemented in the frequency domain, i.e. defining a so-called spectral redundancy rate which quantifies the concept of redundancy among iso-frequency oscillatory components of the analyzed processes. This allows to achieve a non-negative decomposition of the spectral MIR between target and source processes at each specific frequency. Moreover, in the case of linear Gaussian processes Geweke (1982); Chicharro (2011), the use of the minimum MI principle Barrett (2015) applied to the MIR between the target and each source process leads us to retrieve a non-negative time-domain PIRD through integration of the the PI rate atoms over the whole frequency spectrum.
The feasibility of the new framework is displayed exhaustively in benchmark simulations of linearly interacting processes, where we illustrate the benefits of PIRD over the static PID in the presence of non-trivial temporal correlations, investigate the differences with the PID applied to the TE, and highlight the peculiarity of the spectral approach which allows to perform PID restricting the analysis to predetermined frequency bands with practical meaning. The latter property is then exploited in the applicative context of Network Physiology Ivanov (2021) reporting the application of the coarse-grained PIRD to a physiological network comprising cerebrovascular, cardiovascular and respiratory time series analyzed in patients prone to develop postural-related syncope.
@@ II Decomposition of multivariate information
Let us consider a static network system composed of $M+1$ nodes $\mathcal{V}=\{\mathcal{T},\mathcal{S}_{1},\ldots,\mathcal{S}_{M}\}$, where the activity at each node is described in terms of random variables $T,S_{1},\ldots,S_{M}$, with $T$ assumed as the target variable and $S=\{S_{1},\ldots,S_{M}\}$ as the vector of sources. The identification of a target and several source variables naturally leads to use directed approaches to assess and decompose the multivariate information shared in the system, i.e., the mutual information (MI) between $T$ and $S$ defined as
$I(T;S)=\mathbb{E}\left[\log\frac{p(t,s)}{p(t)p(s)}\right],$ (1)
where $p(\cdot,\cdot)$ and $p(\cdot)$ denote joint and marginal probability, with $\mathbb{E}[\cdot]$ being the expectation operator. Directed approaches to information decomposition expand the MI in (1) into terms related to the contributions that the individual sources $S_{i}$, $i=1,\ldots,M$, or a collection thereof, share with the target $T$. Such contributions are related in non-trivial ways to the marginal MI between each source and the target, $I(T;S_{i})$. In this context, the partial information decomposition (PID) introduced by Williams and Beer (Williams and Beer (2010)) provides one of the most popular approaches to study multivariate systems through the lens of a directed decomposition of multivariate information.
@@ II.1 Partial Information Decomposition and the redundancy lattice
Considering the target variable $T$ and the vector source variable $S$, the mathematical redundancy lattice structure defined for the PID Williams and Beer (2010) identifies a set of atoms whose associated partial information (PI) amounts constitute the building blocks of the analyzed multivariate information $I(T;S)$. The lattice is identified by the collection $\mathcal{A}$ of all subsets of sources such that no source is a superset of any other, i.e., the set of anti-chains formed from the indices of the sources in $S$ under set inclusion Williams and Beer (2010) (see Fig. 1a,b (left) for the cases of $M=2$ and $M=3$ sources; e.g, $\mathcal{A}=\{\{1\}\{2\},\{1\},\{2\},\{12\}\}$ if $M=2$). Formally, the PID expands (1) as
$I(T;S)=\sum_{\alpha\in\mathcal{A}}I^{\delta}(T;S_{\alpha}),$ (2)
where $I^{\delta}(\cdot;\cdot)$ is the PI function defined over the atoms of the lattice, $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}\in\mathcal{A}$, and $S_{\alpha}$ denotes the set of subsets of source variables indexed by the $\alpha^{\mathrm{th}}$ atom, with $\alpha_{j}\subseteq\{1\cdots M\}$, $S_{\alpha_{j}}\subseteq S$, $j=1,\ldots,J$.
@@
$I(T;S_{i})=\sum_{\beta\preceq\{i\}}I^{\delta}(T;S_{\beta}),$ (3)
@@
$I^{\cap}(T;S_{\alpha})=\sum_{\beta\preceq\alpha}I^{\delta}(T;S_{\beta}),$ (4)
@@
$I^{\delta}(T;S_{\alpha})=I^{\cap}(T;S_{\alpha})-\sum_{\beta\prec\alpha}I^{\delta}(T;S_{\beta}),$ (5)
or in a compact way via Möbius inversion of (4) Williams and Beer (2010). Importantly, while the redundancy value of an atom $\alpha$, $I^{\cap}(T;S_{\alpha})$, measures the total amount of redundant information shared by all the sources included in that atom, the PI measures the unique information contributed only by that atom.
@@
$I^{\cap}(T;S_{\alpha})=\min\limits_{j=1,\ldots,J}I(T;S_{\alpha_{j}});$ (6)
for instance, considering $M=3$ sources and the atom $\alpha=\{\{1\},\{23\}\}$, the redundancy becomes the minimum between $I(T;S_{1})$ and $I(T;S_{2},S_{3})$. The MMI PID is generally applied in case of Gaussian data, for which it subsumes several of the previously proposed PID schemes Barrett (2015); we will adopt this approach in the practical implementation of the redundancy rate for random processes defined in Sect. "Decomposition of Multivariate Information".
Figure 1: Standard and coarse-grained partial information decomposition, superimposed on the redundancy lattices for $M$ sources. (a) Standard PID on the redundancy lattice for 3 variables ($M=2$). The alphabet of source combinations is $\mathcal{A}=\{\{1\}\{2\},\{1\},\{2\},\{12\}\}$, where $\{1\}$ denotes $S_{1}$ and $\{2\}$ denotes $S_{2}$. The MI between the target and the set of sources is decomposed into a redundant (magenta), a synergistic (blue) and two unique (gray) contributions, which are exclusively provided by different atoms of the lattice. (b) Standard and coarse-grained $1^{st}$ order PID on the redundancy lattice for 4 variables ($M=3$). The alphabet of source combinations is $\mathcal{A}=\{\{1\}\{2\}\{3\},\{1\}\{2\},\{1\}\{3\},\{2\}\{3\},\ldots,\{123\}\}$, where $\{i\}$ denotes $S_{i}$, $i=1,\ldots,M$. The MI between the target and the set of sources is decomposed into a redundant (magenta), a synergistic (blue) and three unique (gray) contributions, which are exclusively provided by different set of atoms of the lattice.

### v3

@@ Abstract
The Partial Information Decomposition (PID) framework has emerged as a powerful tool for analyzing high-order interdependencies in complex network systems. However, its application to dynamic processes remains challenging due to the implicit assumption of memorylessness, which often falls in real-world scenarios. In this work, we introduce the framework of Partial Information Rate Decomposition (PIRD) that extends PID to random processes with temporal correlations. By leveraging mutual information rate (MIR) instead of mutual information (MI), our approach decomposes the dynamic information shared by multivariate random processes into unique, redundant, and synergistic contributions obtained aggregating information rate atoms in a principled manner. To concretely implement this idea, we define a pointwise redundancy rate function based on the minimum MI principle applied locally in the frequency-domain representation of the processes. The framework is validated in benchmark simulations of Gaussian systems, demonstrating its advantages over traditional PID in capturing temporal correlations and showing how the spectral representation may reveal scale-specific higher-order interactions that are obscured in the time domain. Furthermore, we apply PIRD to a physiological network comprising cerebrovascular and cardiovascular variables, revealing frequency-dependent redundant information exchange during a protocol of postural stress. Our results highlight the necessity of accounting for the full temporal statistical structure and spectral content of vector random processes to meaningfully perform information decomposition in network systems with dynamic behavior such as those typically encountered in neuroscience and physiology.
@@ I Introduction
Thanks to the peculiarities described above, the PID framework has become widely adopted as a main tool to assess high-order interdependencies among the units of network datasets collected in several applicative fields of physics, engineering and life sciences where understanding the specific contributions of individual variables to the collective information is of fundamental importance Wibral et al. (2017); Cang and Nie (2020); Rosas et al. (2020); Luppi et al. (2022); Wollstadt et al. (2023); Dissanayake et al. (2024). For instance, PID has been used to analyze neural information processing dynamics, revealing how different brain regions encode stimuli both redundantly and synergistically Kay et al. (2022). In machine learning, PID has been exploited in feature selection to distinguish between information that is redundant across features and information that is uniquely informative Barnett et al. (2009a); Taylor (2018); Wollstadt et al. (2023); Barà et al. (2024). Recent applications have highlighted its value for network analysis also in computational neuroscience, biology and physiology Luppi et al. (2024a); Sherrill et al. (2021); Faes et al. (2021).
@@
Another crucial issue, which is the main question addressed in this paper, is how to apply PID to random processes with temporal statistical structure. In fact, although specifically defined for random variables, PID is often needed in practice to analyze multivariate time series whose most proper statistical representation is the vector random process. Then, since both the target random process and the set of source processes constitute collections of random variables, utilization of the PID in such dynamic case is not straightforward as it implies an arbitrary selection of the variables to be extracted from each process. In the literature, PID has been applied to the variables sampling the processes at the same time, performing a so-called "static PID", under the implicit assumption that the processes are stationary and memoryless (i.e., composed by independent and identically distributed - i.i.d. - variables). However, the i.i.d. assumption is typically not tested in practice and is often violated in applications of information decomposition where the analyzed data exhibit nontrivial temporal correlations Kay et al. (2022); Varley et al. (2023a, b). Alternatively, the PID has been applied to random processes by selecting the variables to use in order to decompose the joint information transferred from all sources to the target Krohova et al. (2019); Luppi et al. (2022); Koçillari et al. (2023); Luppi et al. (2024b). Specifically, taking the present of one predefined process as the target variable and the past histories of all other processes as (vector) source variables, and conditioning the MI between target and sources on the past of the predefined process, yields the PID of the popular transfer entropy (TE) measure of information transfer Schreiber (2000). Nevertheless, although it considers the temporal statistical structure of the multivariate process, the PID of the TE cannot account for instantaneous interactions among the processes, nor for causal interactions occurring in the causal direction from the target to the sources. Hence, the current applications of PID to random processes provide only a partial, often misleading view of the interactions among the units of dynamic network systems.
@@
To deal with the significant issues outlined above, the present work introduces a framework for the decomposition of the information shared dynamically between the target and the source units composing the analyzed network of random processes. Our idea is to shift the paradigm of PID from the use of random variables to the use of random processes as building blocks of information decomposition: leveraging information rate quantities in place of standard information quantities, we replace the MI between random variables with the MI rate (MIR) between random processes, and use the same lattice backbone of PID to implement the so-called partial information rate decomposition (PIRD). PIRD dissects the information shared per unit of time between the designed target random process and the set of source processes, quantified by the MIR, into PI rate atoms adopting a full PID perspective, or into unique, redundant and synergistic information rates adopting a coarse-graining perspective. The PIRD is then completed by introducing the new information-theoretic measure of redundancy rate which generalizes the MIR over the lattice. Among several possible definitions, the redundancy rate is here formulated for the first time following a pointwise approach implemented in the frequency domain, i.e. quantifying the concept of redundancy among iso-frequency oscillatory components of the analyzed processes. Specifically, the redundancy rate is defined first at each specific frequency applying the minimum MI principle Barrett (2015) to the spectral MIR between the target and source processes, and then in the time domain by integrating over all frequencies. This allows to retrieve a non-negative decomposition of the MIR, both locally in frequency and globally in the time domain. Furthermore, a computationally reliable implementation of the framework is provided for the case of linear Gaussian random processes for which the spectral MIR can be defined from the power spectral density matrix Geweke (1982); Chicharro (2011). The feasibility of this implementation is tested exhaustively in benchmark simulations of linearly interacting processes, where we illustrate the benefits of PIRD over the static PID in the presence of non-trivial temporal correlations, investigate the differences with the PID applied to the TE, and highlight the peculiarity of the spectral approach which allows to perform PID restricting the analysis to predetermined frequency bands with practical meaning. The latter property is then exploited in the applicative context of Network Physiology Ivanov (2021) reporting the application of the coarse-grained PIRD to a physiological network comprising cerebrovascular, cardiovascular and respiratory time series analyzed in patients prone to develop postural-related syncope.
@@ II Decomposition of multivariate information
Let us consider a static network system composed of $M+1$ nodes $\mathcal{V}=\{\mathcal{T},\mathcal{S}_{1},\ldots,\mathcal{S}_{M}\}$, where the activity at each node is described in terms of random variables $T,S_{1},\ldots,S_{M}$, with $T$ assumed as the target variable and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}=\{S_{1},\ldots,S_{M}\}$ as the vector of sources. The identification of a target and several source variables naturally leads to use directed approaches to assess and decompose the multivariate information shared in the system, i.e., the mutual information (MI) between $T$ and S defined as
$I(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})=\mathbb{E}\left[\log\frac{p(t,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{s}})}{p(t)p({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{s}})}\right],$ (1)
where $p(\cdot,\cdot)$ and $p(\cdot)$ denote joint and marginal probability, with $\mathbb{E}[\cdot]$ being the expectation operator. Directed approaches to information decomposition expand the MI in (1) into terms related to the contributions that the individual sources $S_{i}$, $i=1,\ldots,M$, or a collection thereof, share with the target $T$. Such contributions are related in non-trivial ways to the marginal MI between each source and the target, $I(T;S_{i})$. In this context, the partial information decomposition (PID) introduced by Williams and Beer Williams and Beer (2010) provides one of the most popular approaches to study multivariate systems through the lens of a directed decomposition of multivariate information.
@@ II.1 Partial Information Decomposition and the redundancy lattice
Considering the target variable $T$ and the vector source variable S, the mathematical redundancy lattice structure defined for the PID Williams and Beer (2010) identifies a set of atoms whose associated partial information (PI) amounts constitute the building blocks of the analyzed multivariate information $I(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})$. The lattice is identified by the collection $\mathcal{A}$ of all subsets of sources such that no source is a superset of any other, i.e., the set of anti-chains formed from the indices of the sources in S under set inclusion Williams and Beer (2010) (see Fig. 1a,b (left) for the cases of $M=2$ and $M=3$ sources; e.g, $\mathcal{A}=\{\{1\}\{2\},\{1\},\{2\},\{12\}\}$ if $M=2$). Formally, the PID expands (1) as
$I(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}})=\sum_{\alpha\in\mathcal{A}}I^{\delta}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\alpha}),$ (2)
where $I^{\delta}(\cdot;\cdot)$ is the PI function defined over the atoms of the lattice, $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}\in\mathcal{A}$, and ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\alpha}$ denotes the set of subsets of source variables indexed by the $\alpha^{\mathrm{th}}$ atom, with $\alpha_{j}\subseteq\{1\cdots M\}$, $_{\alpha_{j}}\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}$ S, $j=1,\ldots,J$.
Figure 1: Standard and coarse-grained partial information decomposition, superimposed on the redundancy lattices for $M$ sources. (a) Standard PID on the redundancy lattice for 3 variables ($M=2$). The alphabet of source combinations is $\mathcal{A}=\{\{1\}\{2\},\{1\},\{2\},\{12\}\}$, where $\{1\}$ denotes $S_{1}$ and $\{2\}$ denotes $S_{2}$. The MI between the target and the set of sources is decomposed into a redundant (magenta), a synergistic (blue) and two unique (grey) contributions, which are exclusively provided by different atoms of the lattice. (b) Standard and coarse-grained $1^{st}$ order PID on the redundancy lattice for 4 variables ($M=3$). The alphabet of source combinations is $\mathcal{A}=\{\{1\}\{2\}\{3\},\{1\}\{2\},\{1\}\{3\},\{2\}\{3\},\ldots,\{123\}\}$, where $\{i\}$ denotes $S_{i}$, $i=1,\ldots,M$. The MI between the target and the set of sources is decomposed into a redundant (magenta), a synergistic (blue) and three unique (grey) contributions, which are exclusively provided by different set of atoms of the lattice.
@@
$I(T;S_{i})=\sum_{\beta\preceq\{i\}}I^{\delta}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\beta}),$ (3)
@@
$I^{\cap}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\alpha})=\sum_{\beta\preceq\alpha}I^{\delta}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\beta}),$ (4)
@@
$I^{\delta}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\alpha})=I^{\cap}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\alpha})-\sum_{\beta\prec\alpha}I^{\delta}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\beta}),$ (5)
or in a compact way via Möbius inversion of (4) Williams and Beer (2010). Importantly, while the redundancy value of an atom $\alpha$, $I^{\cap}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\alpha})$, measures the total amount of redundant information shared by all the sources included in that atom, the PI measures the unique information contributed only by that atom.
@@
$I^{\cap}(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\alpha}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=I^{\mathrm{MMI}}(T;\textbf{S}_{\alpha}):=}\min\limits_{j=1,\ldots,J}I(T;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{S}}_{\alpha_{j}});$ (6)
for instance, considering $M=3$ sources and the atom $\alpha=\{\{1\},\{23\}\}$, the redundancy becomes the minimum between $I(T;S_{1})$ and $I(T;S_{2},S_{3})$. Although it tends to overestimate the amount of redundancy expected by intuition Timme et al. (2014); Griffith and Ho (2015); Harder et al. (2013), the MMI PID is widely used thanks to its simplicity and generality; in particular, it is generally applied in case of Gaussian data, for which it subsumes several of the previously proposed PID schemes Barrett (2015). We will adopt a modified version of this approach in the practical implementation of the redundancy rate for random processes defined in Sect. "Decomposition of Multivariate Information".